OpenStudy (anonymous):

Given the system of equations, match the following items. x + 3y = 5 x - 3y = -1 {5 3} {-1 -3} {1 5} {1 -1} {1 3} {1 -3} is the first set an x, y or system determinant?

OpenStudy (anonymous):

write slope intercept ?

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

x + 3y = 5 time (-1) then : -x-3y =-5 -x-3y =-5 x-3y = -1 ADD------------ -6y =-6 y=-6/-6 y =1 Subtitution y=1 solve for x x + 3(1) = 5 x +3 =5 x =5-3 x =2

OpenStudy (anonymous):

I don't see match with your answer,check your question